Give an example for each of the following:


A set having no limit point.

1
Expert's answer
2022-03-13T18:07:59-0400

Definition. A limit point (accumulation point) of a set "S" in a metric space "X" is a point "x\in X" any puncted neighborhood of which contains an element of "S". The set of all limit points of "S" is denoted by "S'".


In other words, "S'" is the set of all points "x\in X" at which the set "S" is locally infinite. Really, if there exists a neighborhood "U" of "x" such that "U\cap S" is finite then, by putting "r=\min\{dist(x,y): y\in U\cap S,y\ne x\}", we have that a puncted open ball "\dot{B}_r(x)" contains no elements of "S". This means that "x" is not a limit point of "S". Conversely, if any neighborhood "U" of "x" contains an infinite subset of "S" then a puncted neighborhood "\dot{U}" of "x" also contains an infinite subset of "S", hence, "x" is a limit point of "S".


Examples of locally finite sets (i.e. sets having no limit point):

1) any finite set;

2) the set of integers "\mathbb{Z}\subset\mathbb{R}" (any open interval of length "l" contains "[l]" or less integers);

3) the set "S=\{\frac{1}{n}:n\in\mathbb{N}\}" in the metric space "X=(0,+\infty)" with the standard metric. Indeed, if "x=\frac{1}{n}\in S" then the interval "U=(\frac{1}{n+1}, \frac{1}{n-1})" is a neighborhood of "x" such that the puncted "U" contains no elements of "S". Therefore, "x\notin S'". If "x\notin S" and "x>0" then put "n=[1\/x]".We have "n<1\/x<n+1", thus, "\frac{1}{n+1}<x<\frac{1}{n}". The interval "U=(\frac{1}{n+1}, \frac{1}{n})" is an open neighborhood of "x" containing no elements of "S". Thus, "x\notin S'" and therefore, "S'=\emptyset".